OpenStudy (routohan75):

Transform this equation to slope-intercept form: 6x + 4y = 8. A. y=-3/2x+4 B. y=-3/2x+2 C. y=3/2x+4 D. y=3/4x+2

OpenStudy (routohan75):

I have no clue how to transform the equation.

OpenStudy (andrewishere):

So its 6x+4y=8 This is in the form mx+y=b We need it in the form y=mx+b So b, which is 8 is fine because it is on the correct side. So move the mx, which is 6x, to the right side of the equation by undoing it, which would make it negative (-6x) Now the equation looks like this: 4y=-6x+8 But we need y by itself so... we are going to divide everything by 4, which will isolate y. So that looks like this, -6/4 reduces to the fraction -2/3 and 8/4=2. So the final equation is y=mx+b format looks like this y=-2/3x+2 Hopw this helped